OpenStudy (anonymous):

Write an equation in slope intercept form that is parallel to y=3/4x+2 and goes through (8,1)

OpenStudy (isaiah.feynman):

Everyone is asking this type of question.

OpenStudy (anonymous):

Do you know how because I don't

OpenStudy (anonymous):

The slope of the parallel line is going to have the same exact slope. So 3/4 Plug in everything into the point slope equation. y - y1 = m(x - x1) y - 1 = (3/4)(x - 8) y = (3/4)(x-8) + 1 Now just simplify the equation to slope intercept form. y = (3/4)(x-8) + 1 y = (3/4)x - 6 + 1 y = (3/4)x - 5

OpenStudy (anonymous):

Thanks a lot!!!